OpenStudy (anonymous):

Is the sequence –9.9, –13, –16, –19.2, … arithmetic? If so, identify the common difference.

OpenStudy (anonymous):

What's the change between -9.9 and -13?

OpenStudy (anonymous):

-22.9

OpenStudy (anonymous):

no wait, its -3.1

OpenStudy (amistre64):

does: a1-a2 = a2 - a3 ? or, a1+a3 = 2 a2 or, (a1+a3)/2 = a2? if so, then its bound to be arithmetic

OpenStudy (amistre64):

as is, i dont see how that 16+9.9 can half out to 13

OpenStudy (anonymous):

A. yes, –3.2 B. yes, –2.9 C. yes, –3.1 D. no

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

do you see a way that (16+9.9)/2 = 13 ?? if not, then I know which one id pick